OpenStudy (anonymous):

solve for x. 3ln(x-2)=-6

OpenStudy (zzr0ck3r):

divide both sides by 3 and what do you have?

OpenStudy (anonymous):

ln(x-2)=-2

OpenStudy (zzr0ck3r):

now raise both sides to the power of e \(e^{ln(x-2)}=e^{-2}\) the e and ln "cancel" on the left hand side \(x-2=e^{-2}\)

OpenStudy (zzr0ck3r):

solve for x

OpenStudy (anonymous):

\[x=e ^{-2} +2\]

OpenStudy (zzr0ck3r):

correct